OpenStudy (anonymous):

3. A drama club is planning a bus trip to New York City to see a Broadway play. The cost per person for the bus rental varies inversely as the number of people going on the trip. It will cost $30 per person if 44 people go on the trip. How much will it cost per person if 60 people go on the trip? Round your answer to the nearest cent, if necessary. $22.00 $40.91 $1,320.00 $21.29

OpenStudy (kinggeorge):

Again, you start with the equation\[y=\frac{k}{x}\]since it's inversely proportional. What should we now put in for x and y?

OpenStudy (anonymous):

y=30? and x=60?

OpenStudy (kinggeorge):

We're assuming x is the number of people, and y is the cost per person. We know that when 44 people go, it costs 30 dollars per person. So we use y=30 and x=44.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so it would be 1,320.00

OpenStudy (kinggeorge):

That's your \(k\), but not the price per person. Now you need to go back to \[y=\frac{k}{x}\] except now you know k, and you're told to find y when x=60.

OpenStudy (anonymous):

u are smart with math lol u wanna do all my math work lol jk jk jk

OpenStudy (anonymous):

1. Suppose that y varies jointly with w and x and inversely with z and y = 540 when w = 15, x = 30, and z = 5. Write the equation that models the relationship. (1 point) (0 pts) y = (0 pts) y = (1 pt) y = (0 pts) y = 1 /1 point 2. A drama club is planning a bus trip to New York City to see a Broadway play. The cost per person for the bus rental varies inversely to the number of people going on the trip. It will cost $30 per person if 44 people go on the trip. How much will it cost per person if 60 people go on the trip? Round your answer to the nearest cent, if necessary. (1 point) (1 pt) $22.00 (0 pts) $40.91 (0 pts) $1,320.00 (0 pts) $21.29 1 /1 point 3. Graph the function. xy + 16 = 0 (1 point) (1 pt) 3/6/2015 Rational Functions Unit Test https://www.connexus.com/assessments/engine.aspx?idWebuserAssessment=216008011&mode=review&resize=true&popup=true&close=true 2/14 (0 pts) (0 pts) (0 pts) 3/6/2015 Rational Functions Unit Test https://www.connexus.com/assessments/engine.aspx?idWebuserAssessment=216008011&mode=review&resize=true&popup=true&close=true 3/14 1 /1 point 4. Sketch the asymptotes and graph the function (1 point) (1 pt) 3/6/2015 Rational Functions Unit Test https://www.connexus.com/assessments/engine.aspx?idWebuserAssessment=216008011&mode=review&resize=true&popup=true&close=true 4/14 (0 pts) (0 pts) (0 pts) 1 /1 point 5. What is the equation of the function y = translated 3 units to the left and 4 units up? (1 point) 3/6/2015 Rational Functions Unit Test https://www.connexus.com/assessments/engine.aspx?idWebuserAssessment=216008011&mode=review&resize=true&popup=true&close=true 5/14 (0 pts) (0 pts) (0 pts) (1 pt) 0 /1 point 6. Find any points of discontinuity for the rational function. y = (1 point) (0 pts) x = 1, x = 7 (0 pts) x = 8 (1 pt) x = 1, x = –7 (0 pts) x = –1, x = 7 1 /1 point 7. Describe the vertical asymptote(s) and hole(s) for the graph of y = . (1 point) (0 pts) asymptotes: x = –2, 1 and no holes. (0 pts) asymptotes: x = 2, –1 and hole: x = 4. (1 pt) asymptotes: x = –2, –1 and no holes. (0 pts) asymptotes: x = –2, –1 and hole: x = 4 0 /1 point 8. Find the horizontal asymptote of the graph of y = (1 point) (1 pt) y = – (0 pts) y = 0 (0 pts) y = 1 (0 pts) no horizontal asymptote 1 /1 point 3/6/2015 Rational Functions Unit Test https://www.connexus.com/assessments/engine.aspx?idWebuserAssessment=216008011&mode=review&resize=true&popup=true&close=true 6/14 9. What is the graph of the rational function? y = (1 point) (0 pts) (1 pt) 3/6/2015 Rational Functions Unit Test https://www.connexus.com/assessments/engine.aspx?idWebuserAssessment=216008011&mode=review&resize=true&popup=true&close=true 7/14 (0 pts) (0 pts) 0 /1 point 10. Simplify the rational expression. State any restrictions on the variable. (1 point) 3/6/2015 Rational Functions Unit Test https://www.connexus.com/assessments/engine.aspx?idWebuserAssessment=216008011&mode=review&resize=true&popup=true&close=true 8/14 (1 pt) , t ≠ 3, t ≠ –3 (0 pts) , t ≠ 3 (0 pts) , t ≠ 3, t ≠ –3 (0 pts) , t ≠ 3, t ≠ –2 1 /1 point 11. What is the product in simplest form? State any restrictions on the variable. • (1 point) (0 pts) , x ≠ –3, x ≠ –4 (1 pt) , x ≠ –3, x ≠ –4, x ≠ 1 (0 pts) , x ≠ –3, x ≠ –4, x ≠ 1 (0 pts) , x ≠ –3, x ≠ –4 1 /1 point 12. Which is equivalent to ? (1 point) (0 pts) (0 pts) (1 pt) (0 pts) 0 /1 point 13. What is the difference in simplest form? – 3/6/2015 Rational Functions Unit Test https://www.connexus.com/assessments/engine.aspx?idWebuserAssessment=216008011&mode=review&resize=true&popup=true&close=true 9/14 (1 point) (1 pt) (0 pts) (0 pts) (0 pts) 0 /1 point 14. Simplify the sum. (1 point) (0 pts) (0 pts) (0 pts) (1 pt) 0 /1 point 15. Simplify the complex fraction. (1 point) (1 pt) (0 pts) (0 pts) (0 pts) 3/6/2015 Rational Functions Unit Test https://www.connexus.com/assessments/engine.aspx?idWebuserAssessment=216008011&mode=review&resize=true&popup=true&close=true 10/14 1 /1 point 16. Solve the equation. Check the solution. (1 point) (0 pts) –16, –4 (0 pts) 4 (0 pts) –16, 4 (1 pt) –16 1 /1 point 17. A group of college students is volunteering for Help the Homeless during their spring break. They are putting the finishing touches on a house they built. Working alone, Irina can paint a certain room in 8 hours. Paulo can paint the same room in 7 hours. Write an equation that can be used to find how long it will take them working together to paint the room. How many hours will it take them to paint the room? If necessary, round your answer to the nearest hundredth. (1 point) (0 pts) + = 1; 15 hours (0 pts) + = ; 3.73 hours (1 pt) + = 1; 3.73 hours (0 pts) + = 1; 1.87 hours 1 /1 point Short Answer Note: Your teacher will review your response to ensure you receive proper credit for your answer. 18. Determine whether the relation shown in the table is a direct variation, an inverse variation, or neither. If it is a direct or inverse variation, write an equation for the function. (2 points) Essay: Its an inverse variation y=5/x 2 /2 points Nice job finding that it is inverse variation and finding the equation!